# Seemingly unrelated Bayesian additive regression trees for cost-effectiveness analyses in healthcare

### v1

@@ Abstract
In recent years, theoretical results and simulation evidence have shown Bayesian additive regression trees to be a highly-effective method for nonparametric regression. Motivated by cost-effectiveness analyses in health economics, where interest lies in jointly modelling the costs of healthcare treatments and the associated health-related quality of life experienced by a patient, we propose a multivariate extension of BART applicable in regression and classification analyses with several correlated outcome variables. Our framework overcomes some key limitations of existing multivariate BART models by allowing each individual response to be associated with different ensembles of trees, while still handling dependencies between the outcomes. In the case of continuous outcomes, our model is essentially a nonparametric version of seemingly unrelated regression. Likewise, our proposal for binary outcomes is a nonparametric generalisation of the multivariate probit model. We give suggestions for easily interpretable prior distributions, which allow specification of both informative and uninformative priors. We provide detailed discussions of MCMC sampling methods to conduct posterior inference. Our methods are implemented in the R package suBART. We showcase their performance through extensive simulations and an application to an empirical case study from health economics. By also accommodating propensity scores in a manner befitting a causal analysis, we find substantial evidence for a novel trauma care intervention’s cost-effectiveness.
@@ 1 Introduction
Many research questions in health economics are concerned with trading off the costs and benefits of a medical intervention. The most prominent examples are in cost-effectiveness analysis (CEA), where we wish to decide whether a new innovative treatment is worth the associated increase in costs compared to usual care. We therefore need to estimate the average treatment effects on both costs and health. However, in order to get coherent measures of uncertainty, the two treatment effects must be estimated jointly in order to account for the correlation between them (Baio, 2012). This point is elaborated in Section 2. If the CEA is performed with observational data, where the treatment assignment is not randomised, we additionally have to adjust for confounding bias in the analysis.
@@
Chipman et al. (2010) introduced BART as an ensemble method, where each learner is a tree following the Bayesian CART approach previously proposed by the same authors (Chipman et al., 1998). Considering a univariate response vector $\mathbf{y}\in\mathbb{R}^{n}$ and a set of predictors $\mathbf{X}\in\mathbb{R}^{n\times p}$, which may be of mixed type, one of the main objectives of regression modelling is to estimate the conditional expectation $\mathbb{E}[y_{i}\>|\>\mathbf{x}_{i}]=f(\mathbf{x}_{i})$. As a nonparametric model, BART offers high flexibility in estimating this conditional expectation. However, the propensity of decision trees to overfit is mitigated by the Bayesian underpinnings of the method allowing informative prior distributions to impose regularisation in a principled and transparent fashion, as well as the additive nature of the ensemble. These features facilitate better generalisation, in a similar vein to the gradient boosting approach of Friedman (2001). The success of BART is widely reported in the literature across a broad spectrum of applications (Janizadeh et al., 2021; Sarti et al., 2023; Yee and Deshpande, 2023). In addition, theoretical work has demonstrated the frequentist optimality of BART under certain conditions (Linero and Yang, 2018; Ročková and Saha, 2019; Rocková et al., 2020; Rocková, 2020).
@@
The BART model was originally developed for univariate responses but, naturally, subsequent formulations emerged to adapt BART to multivariate responses. Examples include the Bayesian additive vector autoregressive tree (Huber and Rossini, 2022) for multivariate time-series analysis and formulations by Peruzzi and Dunson (2022) for multivariate spatial data. Other applied work involves an extension of BART to forecast the tails of multivariate responses (Clark et al., 2023). Um et al. (2023) adapted BART to cover not only multivariate responses but also the assumption of a skewed-normal distribution. Furthermore, McJames et al. (2023) proposed an extension of Bayesian causal forests (BCF; Hahn et al., 2020) for multivariate responses. All of the aforementioned approaches share the limitation that the tree structure must be identical for each component of the outcome vector. It is easy to envisage situations for which this approach is inappropriate: a specific covariate may be strongly associated with one outcome but independent of another. For example, factors which govern costs may be unrelated to quality of life, and vice versa.
@@
In this study, we propose a novel variant of BART termed seemingly unrelated BART (suBART) which is designed to handle multivariate continuous responses and address this key limitation. Chipman et al. (2010) previously drew parallels to SUR models (Zellner, 1962) and alluded to the potential extension of BART in this direction. Our framework differs from the aforementioned multivariate BART extensions, which assume a single set of trees with correlated multivariate Gaussian distributions in the terminal nodes. Instead, we jointly fit individual ensembles of trees and model the interdependence of the outcomes through correlated error terms. Thus, suBART also differs from merely applying entirely separate univariate BART models to each outcome. Motivated by our investigation into the cost-effectiveness of the TTCM intervention, we further extend suBART to incorporate propensity scores, in the spirit of Hahn et al. (2020), as befits causal analyses. Beyond CEA settings, we also develop probit suBART, an extension of suBART to accommodate multivariate binary outcomes. We envision this version of the model being useful in economic applications with correlated binary outcomes; see Ramful and Zhao (2009) as an example. Our approach is similar to that of Chakraborty (2016), who presented a version of seemingly unrelated BART for exclusively continuous outcomes with an adaptive number of trees. However, it is not specifically tailored to causal inference objectives typical of CEAs and lacks an available open-source software implementation. We address these gaps by providing a comprehensive framework, which covers either continuous or binary outcomes, and a practical implementation through an R package named suBART [^1] Available at: https://github.com/MateusMaiaDS/suBART..
@@
This article proceeds as follows: Section 2 provides background theory on CEA and motivates the development of the suBART model in the context of the application to the TTCM data from Wiertsema et al. (2019). Section 3 then elaborates on the theoretical underpinnings of the suBART methodology and Section 4 discusses the posterior inference for both the multivariate continuous and multivariate binary outcome settings. Section 5 considers different simulation scenarios and discusses the performance of both suBART models compared with standard competitors. The empirical findings of our application of suBART to the TTCM data are presented in Section 6. Finally, Section 7 summarises the proposed methodologies, highlighting both their limitations and potential for further extension, and presents conclusions regarding the health economic application. Additional results, comparisons, and findings are included in the Appendix.
@@ 2 CEA and the suBART model
We now provide more detail about the CEA setting which inspired the suBART model and review some relevant ideas from health economics and causal inference. Detailed treatments can be found in Gabrio et al. (2019) and Li et al. (2023). We defer a description of the specific TTCM data to which we apply suBART to Section 6.
A major motivation to develop the seemingly unrelated BART method was its potential applicability in cost-effectiveness analyses of healthcare treatments. Such analyses are usually performed with data from clinical trials, but there is increasing interest in the analysis of observational data, where the treatment assignment is not randomised. Among epidemiologists, there is widespread consensus that simple parametric models often lead to severe bias and that flexible nonparametric models are preferable for the analysis of observational data (Hernán and Robins, 2024). It seems reasonable to assume that this would extend to the setting of cost-effectiveness analysis, where we want to infer two treatment effects simultaneously. There is, however, a lack of statistical methods fit for these purposes. Given that BART has proven to be very useful for causal inference in the univariate setting (Hill, 2011; Dorie et al., 2019; Rudolph et al., 2023), we consider it a promising method for the multivariate cost-effectiveness setting.
The fundamental problem of cost-effectiveness analyses in health economics is to determine which of two competing healthcare treatments — usually, but not always, for the same disease — should be implemented. Often one is both more effective and more expensive than the other, which raises the question whether the increase in health is worth the added expenses. Henceforth, we let $c_{i}$ and $q_{i}$ respectively denote the healthcare costs and the health-related quality of life associated with a patient. We also suppose that there are two different treatments of interest, $t=0$ and $t=1$. Using the usual potential outcomes notation, we let $c_{i}(t)$ denote the costs associated with patient $i$, had they received treatment $t$. We do likewise for $q_{i}(t)$. We furthermore suppose that there is some vector of baseline characteristics $\mathbf{x}_{i}$ which have an effect on both the outcomes $c_{i}$ and $q_{i}$, as well as the treatment indicator $t$.
@@
We now make the assumption of ignorability, which means that conditional on the baseline covariates $\mathbf{x}_{i}$, the treatment $t$ is independent of the potential outcomes $c_{i}(0)$ and $c_{i}(1)$. Under this assumption, we may rewrite our treatment effect as
$\Delta_{c}=\frac{1}{n}\sum^{n}_{i=1}\mathbb{E}\left[c_{i}\>|\>t=1,\mathbf{x}_{i}\right]-\mathbb{E}\left[c_{i}\>|\>t=0,\mathbf{x}_{i}\right].$
It follows that $\Delta_{c}$ is completely specified by the conditional expectations $\mathbb{E}[c_{i}\>|\>t,\mathbf{x}_{i}]$. We proceed in the same manner for $\Delta_{q}$, the MATE on the patient’s quality of life. It is then customary to combine the two treatment effects into a utility function, the incremental net benefit (INB):
@@
The scalar parameter $\lambda$ is called the willingness-to-pay. Roughly speaking, $\lambda$ quantifies how much cost (in the given currency) a decision-maker is willing to trade for a one-unit increase in healthcare-related quality of life for one patient. The decision rule is then simple: if the INB is at most zero, we say that treatment $1$ is not cost-effective, and treatment 00 should be implemented. If the INB is larger than zero, we consider treatment $1$ to be cost-effective and worthy of being implemented.
@@
$=\Phi\left(\frac{\lambda\mathbb{E}\left[\Delta_{q}\right]-\mathbb{E}\left[\Delta_{c}\right]}{\left(\lambda^{2}\text{Var}\left[\Delta_{q}\right]+\text{Var}\left[\Delta_{c}\right]-2\lambda\text{Cov}\left[\Delta_{q},\Delta_{c}\right]\right)^{1/2}}\right).$
Consequently, the probability of cost-effectiveness depends on the covariance of $\Delta_{c}$ and $\Delta_{q}$. Without the normality assumption, this probability can usually not be found explicitly, but the same principle applies nonetheless: the probability of cost-effectiveness depends on the joint distribution of $\Delta_{c}$ and $\Delta_{q}$ (Löthgren and Zethraeus, 2000; Gabrio et al., 2019). It follows that we must model $c$ and $q$ jointly, as modelling them separately would enforce the unrealistic prior belief that the treatment effects $\Delta_{c}$ and $\Delta_{q}$ are independent. This belief is seldom appropriate, since empirical cost and health data are often strongly correlated (Willan et al., 2004).

### v3

@@ Abstract
In recent years, theoretical results and simulation evidence have shown Bayesian additive regression trees to be a highly-effective method for nonparametric regression. Motivated by cost-effectiveness analyses in health economics, where interest lies in jointly modelling the costs of healthcare treatments and the associated health-related quality of life experienced by a patient, we propose a multivariate extension of BART which is applicable in regression analyses with several dependent outcome variables. Our framework allows for continuous or binary outcomes and overcomes some key limitations of existing multivariate BART models by allowing each individual response to be associated with different ensembles of trees, while still handling dependencies between the outcomes. In the case of continuous outcomes, our model is essentially a nonparametric version of seemingly unrelated regression. Likewise, our proposal for binary outcomes is a nonparametric generalisation of the multivariate probit model. We give suggestions for easily interpretable prior distributions, which allow specification of both informative and uninformative priors. We provide detailed discussions of MCMC sampling methods to conduct posterior inference. Our methods are implemented in the R package subart. We showcase their performance through extensive simulation experiments and an application to an empirical case study from health economics. By also accommodating propensity scores in a manner befitting a causal analysis, we find substantial evidence for a novel trauma care intervention’s cost-effectiveness.
@@ 1 Introduction
Many research questions in health economics are concerned with trading off the costs and benefits of a medical intervention. The most prominent examples are in cost-effectiveness analysis (CEA), where we wish to decide whether a new innovative treatment is worth the associated increase in costs compared to usual care. We therefore need to estimate the average treatment effects on both costs and health. However, in order to get coherent measures of uncertainty, the two treatment effects must be estimated jointly in order to account for the correlation between them (Baio, 2012). Furthermore, if the CEA is performed with observational data, where the treatment assignment is not randomised, we additionally have to adjust for confounding bias in the analysis. These points are elaborated in Section 2.
@@
Chipman et al. (2010) introduced BART as an ensemble method, where each learner is a tree following the Bayesian CART approach previously proposed by the same authors (Chipman et al., 1998). Considering a univariate response vector $\mathbf{y}\in\mathbb{R}^{n}$ and a set of predictors $\mathbf{X}\in\mathbb{R}^{n\times p}$, which may be of mixed type, the main objective of regression modelling is to estimate the conditional expectation $\mathbb{E}[y_{i}\>|\>\mathbf{x}_{i}]$, also called the regression function. As a nonparametric model, BART offers high flexibility in estimating this conditional expectation. The tendency of decision trees to overfit is mitigated by the Bayesian underpinnings of the method, allowing informative prior distributions to impose regularisation in a principled and transparent fashion, as well as the additive nature of the ensemble. These features facilitate better generalisation, in a similar vein to the gradient boosting approach of Friedman (2001). The success of BART is widely reported in the literature across a broad spectrum of applications (Janizadeh et al., 2021; Sarti et al., 2023; Yee and Deshpande, 2023). In addition, theoretical work has demonstrated the frequentist optimality of BART under certain conditions (Linero and Yang, 2018; Rocková, 2020; Rocková and Van der Pas, 2020).
@@
The BART model was originally developed for univariate responses but subsequent formulations emerged to adapt BART to multivariate responses. Examples include the Bayesian additive vector autoregressive tree (Huber and Rossini, 2022) for multivariate time-series and formulations by Peruzzi and Dunson (2022) for multivariate spatial data. Other applied work involves an extension of BART to forecast the tails of multivariate responses (Clark et al., 2023). Um et al. (2023) adapted BART to cover not only multivariate responses but also the assumption of a skew-normal distribution; throughout this paper, we refer to the variant without skewness as mvBART. Furthermore, McJames et al. (2024) proposed an extension of Bayesian causal forests (BCF; Hahn et al., 2020) for multivariate responses. All aforementioned approaches share the limitation that the tree structure must be identical for each component of the outcome vector. It is easy to envisage situations for which this is inappropriate: a specific covariate may be strongly associated with one outcome but independent of another; e.g., factors which govern costs may be unrelated to quality of life, and vice versa.
@@
In this study, we propose a novel variant of BART termed seemingly unrelated BART (suBART) which is designed to handle multivariate continuous responses and address this key limitation. Chipman et al. (2010) previously drew parallels to SUR models (Zellner, 1962) and alluded to the potential for extending BART in this direction. Our framework differs from the aforementioned multivariate BART extensions, which assume a single set of trees with correlated multivariate terminal node parameters. Instead, we jointly fit individual ensembles of trees for each outcome and model their interdependence through correlated error terms. Thus, suBART also differs from merely applying univariate BART models to each outcome independently. Motivated by our investigation into the cost-effectiveness of the TTCM intervention, we further extend suBART to incorporate propensity scores, in the spirit of Hahn et al. (2020), as befits causal analyses. Beyond CEA settings, we also develop probit suBART, an extension of suBART to accommodate multivariate binary outcomes. While not the main focus of this paper, we envision this version of the model being useful in marketing applications with correlated binary outcomes; see the examples in Rossi et al. (2005). Our approach is similar to that of Chakraborty (2016), who offers a version of seemingly unrelated BART for exclusively continuous outcomes with an adaptive number of trees. However, it is not specifically tailored to causal inference objectives typical of CEA and lacks an available open-source software implementation. Moreover, the non-informative inverse-Wishart prior adopted by Chakraborty (2016) on the covariance matrix of the error terms fails to properly calibrate the prior on the residual variance parameters, which is crucial for regularising the contributions of the trees in the ensemble. We address these gaps by providing a comprehensive framework, which covers either continuous or binary outcomes, with appropriately specified priors, and a practical implementation through an R package named subart [^1] Implemented using Rcpp and available at: https://github.com/MateusMaiaDS/subart..
@@
This article proceeds as follows: Section 2 provides background theory on CEA and motivates the development of the suBART model in the context of the application to the TTCM data from Wiertsema et al. (2019). Section 3 then elaborates on the theoretical underpinnings of the suBART methodology and Section 4 discusses the posterior inference for both the multivariate continuous and multivariate binary outcome settings. Section 5 considers different simulation scenarios to compare the predictive performance of suBART against standard competitors. The empirical findings of our application of suBART to the TTCM data are presented in Section 6, along with an additional simulation experiment based on these data which assesses suBART’s estimation of treatment effects in the CEA setting. Finally, Section 7 summarises the proposed methodologies, highlighting both their limitations and potential for further extension, and presents conclusions regarding the health economic application. Additional results, comparisons, and findings are included in the appendices.
@@ 2 CEA and the suBART model
Our main motivation in developing suBART is its potential applicability in cost-effectiveness analyses of healthcare treatments, particularly in cases where the treatment assignment is not randomised. When estimating treatment effects from such observational data, simple parametric models often lead to severe bias and flexible nonparametric models are preferable (Hernán and Robins, 2024). Naturally, such concerns extend to CEA settings, where we want to estimate two treatment effects simultaneously. However, there is a lack of statistical methods fit for such purposes. Given that BART has proven to be very useful for univariate causal inference tasks (Hill, 2011; Dorie et al., 2019; Rudolph et al., 2023), we consider it a promising basis for multivariate CEA.
We continue by sketching some relevant ideas from health economics and causal inference in Section 2.1, drawing on the detailed treatments found in Gabrio et al. (2019) and Li et al. (2023). Subsequently, we give a detailed presentation of the TTCM data in Section 2.2.
2.1 Cost-effectiveness analyses with observational data
The fundamental problem of cost-effectiveness analyses in health economics is to determine which of two competing healthcare treatments — usually, but not always, for the same disease — should be implemented. Often one is both more effective and more expensive than the other, which raises the question whether the increase in health is worth the added expenses. Henceforth, we let $c_{i}$ and $q_{i}$ respectively denote the healthcare costs and the health-related quality of life associated with a patient. We also suppose that there are two different treatments of interest, $t=0$ and $t=1$. Using the usual potential outcomes notation, we let $c_{i}(t)$ denote the costs associated with patient $i$, had they received treatment $t$. We do likewise for $q_{i}(t)$. We furthermore suppose that there is some vector of baseline characteristics $\mathbf{x}_{i}$ which have an effect on both the outcomes $c_{i}$ and $q_{i}$, as well as the treatment indicator $t_{i}$.
@@ 2.1 Cost-effectiveness analyses with observational data
We now make the assumption of ignorability, which means that conditional on the baseline covariates $\mathbf{x}_{i}$, the treatment $t_{i}$ is independent of the potential outcomes $c_{i}(0)$ and $c_{i}(1)$. Under this assumption, we may rewrite our treatment effect as
$\Delta_{c}=\frac{1}{n}\sum^{n}_{i=1}\mathbb{E}\left[c_{i}\>|\>t_{i}=1,\mathbf{x}_{i}\right]-\mathbb{E}\left[c_{i}\>|\>t_{i}=0,\mathbf{x}_{i}\right].$
Thus, $\Delta_{c}$ is completely specified by the conditional expectations $\mathbb{E}[c_{i}\>|\>t_{i},\mathbf{x}_{i}]$. We proceed in the same manner for $\Delta_{q}$, the MATE on the patient’s quality of life. It is then customary to combine the two treatment effects into a utility function, the incremental net benefit (INB):
@@
The scalar parameter $\lambda$ is called the willingness-to-pay. Roughly speaking, $\lambda$ quantifies how much cost (in the given currency) a decision-maker is willing to trade for a one-unit increase in healthcare-related quality of life for one patient. The decision rule is then simple: if the INB is at most zero, we say that treatment $1$ is not cost-effective, and treatment 00 should be implemented. If the INB is larger than zero, we consider treatment $1$ to be cost-effective and worthy of being implemented. Throughout this paper, we give $\lambda$ values in units of €1,000.
@@
$=\Phi\left(\frac{\lambda\mathbb{E}\left[\Delta_{q}\right]-\mathbb{E}\left[\Delta_{c}\right]}{\left(\lambda^{2}\text{Var}\left[\Delta_{q}\right]+\text{Var}\left[\Delta_{c}\right]-2\lambda\text{Cov}\left[\Delta_{q},\Delta_{c}\right]\right)^{1/2}}\right).$ (2)
Consequently, the probability of cost-effectiveness depends on the covariance of $\Delta_{c}$ and $\Delta_{q}$. Without the normality assumption, this probability can usually not be found explicitly, but the same principle applies nonetheless: the probability of cost-effectiveness depends on the joint distribution of $\Delta_{c}$ and $\Delta_{q}$ (Löthgren and Zethraeus, 2000; Gabrio et al., 2019). It follows that we must model $c_{i}$ and $q_{i}$ jointly, as modelling them separately would enforce the unrealistic assumption that the treatment effects $\Delta_{c}$ and $\Delta_{q}$ are independent. This belief is seldom appropriate, since empirical cost and health data are often strongly correlated (Willan et al., 2004). We expound on these points through simulation experiments in Appendix A.